OpenStudy (anonymous):

In Fig. 21-22, particles 1 and 2 are fixed in place, but particle 3 is free to move. If the net electrostatic force on particle 3 due to particles 1 and 2 is zero and L23 = 5.72L12, what is the ratio q1/q2?

OpenStudy (anonymous):

where is the fig 21-22 ?

OpenStudy (anonymous):

L23=5.72L12 E on 3 = 0 i didnt see the figure, but suppose all of the particles lying on same axis, say x-axis: \[F_{13}=F_{23}\] \[kQ_1Q_3/r^2_{13}=kQ_2Q_3/r^2_{23}\] \[Q_1/r^2_{13}=Q_2/r^2_{23}\] \[Q_1r^2_{23}=Q_2r^2_{13}\] the ratio of q1/q2: \[Q_1/Q_2=r^2_{13}/r^2_{23}\] \[Q_1/Q_2=(L_{12})/(L_{23})^2\] \[Q_1/Q_2=(L_{12})^2/(5.72L_{12})^2\] \[Q_1/Q_2=(1/5.72)^2\]

OpenStudy (anonymous):

i think my final answer isn't correct, but thats how you solve it.